John is 5 years older than Yaw,if the sum of their ages is 19.How old is John

Found 2 solutions by Shin123, ikleyn:
Answer by Shin123(626)

Let j represent John's age and let y represent Yaw's age. . So John is 12 years old.

Answer by ikleyn(52797)

.

    J + (J-5) = 19

    2J - 5 = 19

    2J     = 19 + 5 = 24

     J              = 24/2 = 12.    ANSWER

Solved.

---------------

In 173 times BETTER THAN any online solver.